Your neighbor Paul has rented a truck with a loading ramp. The ramp is tilted upward at 35 ∘, and Paul is pulling a large crate up the ramp with a rope that angles 30 ∘ above the ramp. Paul pulls with a force of 750 N. (Force is measured in newtons, abbreviated N.)

What is the magnitude of the horizontal component of his force?

What is the magnitude of the vertical component of his force?
Answer #1

Magnitude of the horizontal component of his force = 750*cos (35+30) degree

= 750* cos 65 degree

= 317 N Answer

Magnitude of the vertical component of his force = 750*sin (35+30) degree

= 750* sin 65 degree

= 680 N Answer
Answer #2

We have to resolve the forceon vertical and horizontal axes.